OpenStudy (anonymous):

What is the probability of winning a lottery in which you must choose 6 numbers from the numbers 1 through 13? a) Assuming that order is unimportant. b) Assuming order is important.

OpenStudy (wolf1728):

The formula for combinations (how to choose r things from a larger set of n things) is:

OpenStudy (wolf1728):

So for part "a" we have 13! / 6!*(7!) = 13* 12* 11* 10* 9 * 8* 7! / 6!*7! = 13* 12* 11* 10* 9 * 8 / 6 * 5 * 4 * 3 * 2 = 13* 11* 2 * 3 * 2 = 1,716 and to check that, here's a combination calculator: http://www.1728.org/combinat.htm

OpenStudy (amistre64):

should we assume duplication is not acceptable?